THE NEAR INFRARED SYSTEMS OF NICKEL CHLORIDE

\underline{LEAH O’BRIEN}, TENEIL KELLERMAN, and AMANDA LAMBETH, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The near infrared electronic transitions of NiCl were recorded from 9,000 to 15,000 cm^{−1} with the Fourier transform spectrometer associated with the McMath-Pierce Solar Telescope, Kitt Peak, AZ. Two new electronic states at 9,100 cm^{−1} and 10,500 cm^{−1} were observed for the first time. The NiCl molecules were produced in a King-type carbon tube furnace operating at 1500 C and at a total pressure of approximately 400 torr. The electronic transition energies and vibrational constants for the states involved in the near infrared systems will be presented.